# 4D Radar-Inertial Odometry based on Gaussian Modeling and Multi-Hypothesis Scan Matching

### v1

@@ VII Conclusion
In this paper we introduce a multi-hypothesis Gaussian-based radar-inertial odometry pipeline. Our methods summarize noisy 4D radar scans using freely positioned 3D Gaussians (like 3DGS [8] and unlike NDT [2]), which then support multi-hypothesis scan matching for increased robustness. We show that the fusion of these methods into an EKF produces better quantitative and qualitative results than existing baselines using traditional scan matching algorithms.
Future works include improving the Gaussian modeling and scan-matching algorithms (in particular, exploring optimization algorithms more direct than gradient descent), modeling kinematic restrictions relevant to the robot platform used in order to reduce noise during egovelocity estimation, further refining the pose estimates by means of Pose Graph Optimization (PGO), incorporating a loop closure module, and creating a global Gaussian map of the environment for the purposes of full SLAM.

### v2

@@ VII Conclusion
In this paper we introduce a Gaussian-based radar-inertial odometry pipeline. Our methods summarize noisy 4D radar scans using freely positioned 3D Gaussians (like 3DGS [10] and unlike NDT [2]), which then support multi-hypothesis scan matching for increased robustness. Through experiments in multiple datasets with multiple radar types, we show that these methods produce results comparable to existing registration algorithms, and even manage to outperform them in several sequences.
Future works include improving the Gaussian models with an optimization algorithm more efficient than gradient descent, better initialization and dynamic management of Gaussians (as in 3DGS), as well as incorporating radar-specific information such as intensity values; improving the robustness of the multi-hypothesis particle system with kinematic priors, dynamically adjusting the hyperparameters of the system, improving the quality of egovelocity estimation by modeling kinematic restrictions relevant to the robot platform used, further refining the pose estimates by means of Pose Graph Optimization (PGO), incorporating a loop closure module, and creating a global Gaussian map of the environment for the purposes of full SLAM.